OpenStudy (cutebunny12):

Which equation can be used to find eighteen percent of eight hundred ninety-three? eighteen over eight hundred ninety-three equals x over one hundred eighteen over one hundred equals x over eight hundred ninety-three one hundred over x equals eight hundred ninety-three over eighteen x over eighteen equals one hundred over eight hundred ninety-three

OpenStudy (taramgrant0543664):

Which one do you think it is?

OpenStudy (cutebunny12):

a or b

OpenStudy (taramgrant0543664):

Yes it's 18/100=x/893 Which is 0.18=x/893 0.18(293)=x So b would be correct